10=5x-9y
12-2x=3y

Answer by mananth(16946)

5 x -9 y = 10 .............1
Total value
2 x + 3 y = 12 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 3
5 x -9 y = 10
6 x + 9 y = 36
Add the two equations
11 x = 46
/ 11
x = 4 1/6
plug value of x in (1)
5 x -9 y = 10
20 8/9 -9 y = 10
-9 y = 10 -20 8/9
-9 y = -10 8/9
y = 1 2/9